OpenStudy (anonymous):

PLEASE HELP. Determine if each binomial is a perfect square, if so factor, if not explain. 100x^2-81.

OpenStudy (anonymous):

D = b^2 - 4ac D = 0^2 - 4 * 100 * 81 D = -32400 D < 0 so no solution

OpenStudy (anonymous):

@MeganRose

OpenStudy (anonymous):

Ah, I see. Thanks so much.

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

Just say @Mason_Jones573 if you need more help

OpenStudy (anonymous):

Alright I will, thank you. (:

OpenStudy (anonymous):

Your welcome :)